What is the reciprocal of 2+i?
Please solve and explain how to get the anwser

Answer by Alan3354(69443)

What is the reciprocal of 2+i?
------------
It's 1/(2+i)
For some reason, people don't like the "i" in the denominator.
Multiply NUM and DEN by its conjugate.
-----------
= (2-i)/(2+i)*(2-i)
= (2-i)/(4+1)
= (2-i)/5